STRUCTURAL DETERMINATION OF THE LINEAR SILICON-CARBON CHAINS SIC_{4} AND SIC_{6}

\underline{VERNITA GORDON}, Department of Physics and Division of Engineering and Applied Sciences, Harvard University and Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; E. NATHAN, A. J. APPONI, M. C. MCCARTHY, C. A. GOTTLIEB, P. THADDEUS, Division of Engineering and Applied Sciences, Harvard University and Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138.

Silicon-carbon chains and rings are of astronomical interest since SiC, SiC_{2}, rhomboidal SiC_{3}, and linear SiC_{4} have all been detected in the circumstellar shell of the evolved carbon star IRC+10216. We have now detected and characterized in the laboratory the rotational spectra of four new linear silicon carbides from SiC_{5} to SiC_{8}. Molecules were formed by an electrical discharge in a supersonic molecular beam and detected by Fourier-transform microwave spectroscopy. For several of these molecules, lines are sufficiently strong that a number of rare isotopic species have been found as well. Experimental substitution structures for SiC_{4} and SiC_{6} have been determined by fitting the bond lengths to the derived moments of inertia of the isotopic species. Agreement between predicted and observed bond lengths indicates that the bonding in these molecules is cumulenic, i.e. that the bonding is characterized by a series of double bonds.